OpenStudy (anonymous):

estimate the difference of the decimals below by rounding to the nearest whole number . 30.516 - 7.190= would 30.516 be 30.000

OpenStudy (anonymous):

and 7.190 7 .000

OpenStudy (anonymous):

30.5 would be 31

OpenStudy (anonymous):

ok and then what

OpenStudy (anonymous):

31 - 7 = ...

OpenStudy (anonymous):

24

OpenStudy (anonymous):

That's what I would get

OpenStudy (anonymous):

but is our answer complete ?

OpenStudy (anonymous):

24 is a whole number so I would believe so

OpenStudy (anonymous):

so 30.516- 7.190 = 24

OpenStudy (anonymous):

done

OpenStudy (anonymous):

When you are estimating decimal correct.

OpenStudy (anonymous):

Unless you are suppose to subtract then estimate which would be 23. Depends which way you are being taught/the book tells you to do

OpenStudy (anonymous):

so the same method would go for a problem like 53.042 - 3.350

OpenStudy (anonymous):

53 would b 54

OpenStudy (anonymous):

Yes exactly the same, anything with .5 will round up anything with .49 and below will round down.

OpenStudy (anonymous):

53 since it's below 53.5 would round down. if it was 53.49 it would round down.

OpenStudy (anonymous):

Or if it was 53.5 it would be 54

OpenStudy (anonymous):

54-3 ?=51

OpenStudy (anonymous):

Or if it was 53.5 it would be 54

OpenStudy (anonymous):

That one would be 53-3 so it'd be 50.....after your decimal is how you tell to round up or down.

OpenStudy (anonymous):

ok now 6.802 6.184 _6.318=

OpenStudy (anonymous):

idk with this one ?

OpenStudy (anonymous):

What are the numbers again? Sorry comp messed up on me

OpenStudy (anonymous):

6.802 6.184 6.318

OpenStudy (anonymous):

6.8 = 7, 6.1 = 6, 6.3 = 6

OpenStudy (anonymous):

got 6 so is that my answer cause i was thinking i had to do more

OpenStudy (anonymous):

Are you subtracting,adding, anything with those numbers?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

am i adding the 6.1 +